# Towards Anytime Optical Flow Estimation with Event Cameras

### v1

@@ IV-D Time-Dense Optical Flow Evaluation
We perform the same comparison on the EVA-FlowSet, which consists of sequences with curved movements. The results, depicted in Fig. 9, clearly indicate a substantial performance improvement achieved by our approach in comparison to E-RAFT. Since our EVA-FlowSet does not provide ground truth optical flow, both E-RAFT and our model rely on checkpoints trained on the DSEC dataset. Fig. 9 presents a visual representation of the evaluation results comparing our method with E-RAFT in diverse real-world scenarios with varying motion speeds. Specifically, in two fast-moving sequences, our model demonstrates a notably superior intermediate RFWL compared to E-RAFT. This occurs due to the nonlinearity of the optical flow caused by the curved trajectory. Consequently, the interpolated intermediate flow obtained from E-RAFT’s prediction proves notably less accurate compared to the time-dense flow predicted by our model. In the other two scenarios with normal-speed sequences, this model and E-RAFT exhibit comparable RFWL performance. When the movement speed is not fast enough during motion along a curved trajectory, the motion can be approximated as linear for a brief duration.
As depicted in Fig. 8, the qualitative results affirm that our model outperforms E-RAFT in distinguishing motion boundaries, even in scenarios with normal speed. Notably, E-RAFT tends to erroneously extend the optical flow from motion regions to non-motion regions. However, since non-motion regions have few events, any inaccuracies in optical flow estimation will have no impact on the results of motion compensation. Consequently, in this specific situation, the RFWL of E-RAFT is comparable to that of this model. Overall, our proposed EVA-Flow is evidenced to be superior in time-dense flow estimation both numerically and qualitatively.
To assess the effectiveness of our time-dense optical flow estimation framework from an alternative perspective, we employ a different approach for the MVSEC dataset. In Tab. 2, we present the zero-shot results of E-RAFT and our model (trained only on DSEC and directly tested on MVSEC, $bins{=}15$). When evaluating the optical flow estimation on the MVSEC dataset using our model, we utilize the same model checkpoint but employ varying bin numbers based on different $dt$ values. As a result of the divergence in motion speeds between the slow-motion MVSEC dataset and the fast-motion DSEC dataset, we assign $bins{=}4$ for $dt{=}1$ and $bins{=}13$ for $dt{=}4$. This setting is based on the number of bins in our model minus $1$, which equals the number of the time-dense optical flow output. Therefore, under this setting, the frame rate of the optical flow output with $dt{=}4$ is exactly equal to the frame rate of the optical flow output with $dt{=}1$. And when $dt{=}4$ is utilized, it corresponds to an ongoing iteration of the SMR module using $dt{=}1$ time step until the optical flow for the entire duration can be forecasted.
By analyzing the data presented in Tab. 2, it is apparent that our model achieves a lower endpoint error (EPE) of $0.96$ at $dt{=}4$, which is less than four times the EPE at $dt{=}1$ ($0.39$). Interestingly, the results indicate that as we progressively iterate and generate optical flow for longer time intervals, the average per-unit-time optical flow error does not increase, but rather decreases. These findings provide clear evidence for the efficacy of our framework in estimating time-dense optical flow. Moreover, this highlights a key benefit of the adaptable implementation of this framework. It allows for flexible adjustment of the number of prediction bins, enabling a reduction in the event rate discrepancy per bin between training and prediction data, thus achieving higher accuracy in the deployment stage.
IV-E Ablation Study
TABLE 5: Ablation on the bins setting.
Bins Latency/FPS EPE $\downarrow$ AE $\downarrow$ 1PE $\downarrow$ 3PE $\downarrow$
6 $20ms$ / $50Hz$ 0.955 3.29 16.7 3.9
11 $10ms$ / $100Hz$ 0.926 3.34 16.1 3.5
15 $7.1ms$ / $140Hz$ 0.895 3.39 16.1 3.3
21 $5ms$ / $200Hz$ 0.877 3.31 15.9 3.2
31 $3.3ms$ / $300Hz$ 0.901 3.37 17.0 3.2
Unified Voxel Grid. Tab. 4 compares the results of our framework on the DSEC test set using Voxel Grid [17] and our proposed Unified Voxel Grid representation. The EPE of our proposed Unified Voxel Grid is $7.3\%$ lower than that of Voxel Grid. This difference stems from the fact that each event bin in the Unified Voxel Grid adheres to a consistent event representation, while the representation patterns of the first and last bins in the Voxel Grid do not match those of the middle bins (refer to Fig. 3).
Investigate Bins Setting. The design of the proposed framework enables an increase of the optical flow output’s frame rate by increasing the number of bins settings in a $100ms$ sample from the DSEC dataset. It is essential to ensure that each bin has a sufficient number of events to encompass the motions taking place within the scene. In this study, the number of event bins varies from $6$ to $31$. Tab. 5 illustrates a consistent decrease in the final endpoint error (EPE) of event optical flow as the number of bins increases from $6$ to $21$. This trend is attributed to the reduction in complexity during optical flow estimation in each iteration when a higher number of bins is used. This reduction is due to a smaller amount of optical flow that needs to be estimated. However, as the number of event bins continues to increase, individual bins start to contain insufficient event information, resulting in a decline in optical flow accuracy when the number reaches $31$. Our final model ($bins=21$) increases the event output frame rate by 20-fold while maintaining high accuracy.
V Conclusion
@@
In this paper, we propose EVA-Flow, a learning-based model for anytime flow estimation based on event cameras. Leveraging the unified voxel grid representation, our network is able to estimate dense motion fields bin-by-bin between two temporal slices. By employing a novel Spatiotemporal Motion Refinement (SMR) module for implicit warp alignment, the proposed EVA-Flow overcomes the low-temporal-frequency issue of event-based optical flow datasets, achieving high-speed and high-frequency flow estimation beyond low-frequency supervised signals. Compared to the state-of-the-art method E-RAFT, our approach achieves competitive accuracy, fast inference, arbitrary-time flow estimation, and strong generalization. Real-time arbitrary-time flow estimation is achieved on a single GPU, showcasing the significant potential of EVA-Flow for online visual tasks. We also propose an unsupervised event-based optical flow evaluation metric, referred to as Rectified Flow Warp Loss (RFWL), to validate the time-dense optical flow predictions of our model. In the future, we intend to seamlessly integrate EVA-Flow into event-based visual odometry to achieve fast, accurate, and high-dynamic-range pose estimation. Unsupervised domain adaptation for fine-grained visual tasks via event cameras would also be a direction of interest to enhance autonomous perception.

### v3

@@ 4.5 Ablation Study
Number of Bins (N) Setting.
The number of bins ‘N‘ in the UVG representation is a crucial hyperparameter that dictates the temporal granularity ($\tau=T/(N-1)$, where T is the total time window, e.g., 100ms) and the maximum prediction frequency. Setting ‘N‘ involves a trade-off: higher ‘N‘ allows for finer temporal resolution and potentially higher prediction frequency, but requires sufficient event density within each shorter interval $\tau$.
Impact During Training (DSEC): We first investigated the effect of ‘N‘ during training on the DSEC dataset, evaluating the final End-Point Error (EPE). As shown in Tab. 4 (left), increasing ‘N‘ from 6 up to 21 leads to a consistent decrease in the final EPE. This improvement can be attributed to the network estimating smaller displacements in each recurrent step, simplifying the estimation task per iteration. However, when ‘N‘ increases further (e.g., to 31), the accuracy declines as individual bins start to contain insufficient event information, hindering robust flow estimation. Based on these results, we selected $N=21$ for our primary model trained on DSEC, as it provides a good balance, achieving a 20-fold increase in potential output frequency compared to standard methods while maintaining high final accuracy.
Impact During Inference (EVA-FlowSet [$T=20ms$] & Guidance): While training with $N=21$ yields strong results for the final flow prediction, the optimal ‘N‘ for achieving the best time-dense flow quality during inference can vary depending on scene dynamics. To explore this, we conducted experiments on our EVA-FlowSet, evaluating the average time-dense RFWL using different inference values of ‘N‘ (from 2 to 21) on sequences with varying speeds, using the model trained with $N=15$. The results, presented in Fig. 10, reveal that the optimal inference ‘N‘ indeed depends on motion speed. For high-speed sequences, peak RFWL (best time-dense quality) is achieved around $N$=11-12 ($\tau\approx 2$ms), suggesting that fast motion generates enough events for reliable estimation even with short integration times. Conversely, for normal-speed sequences, performance peaks around $N$=9 ($\tau=2.5$ms), indicating that slightly longer integration times are beneficial when motion is slower.
This analysis highlights the flexibility of EVA-Flow. While trained with a specific ‘N‘ (e.g., 21), the inference can be adapted. We provide the following guidance for practical application (for our EVA-Set):
  • For high-speed scenarios, using $\tau\approx 2ms$ during inference is recommended for optimal time-dense performance.
  • For moderate/low-speed or low-light (sparse event) scenarios, using a longer $\tau>2.5ms$ during inference is likely beneficial to ensure sufficient event accumulation per bin.
It is important to note that these ranges serve as initial recommendations. The optimal setting for $\tau$ (and consequently ‘N‘) depends strongly on the specific characteristics of the scene (motion speed, light condition). This adaptability allows users to tune the inference granularity for the specific scenario, maximizing the quality of the time-dense optical flow output.
5 Limitations and Future Work
While EVA-Flow demonstrates significant advantages in low-latency, high-frequency optical flow estimation with high computational efficiency, we acknowledge several limitations that suggest avenues for future research:
  • Accuracy Relative to Non-Time-Dense Methods: Although competitive, EVA-Flow’s final prediction accuracy (e.g., EPE) is slightly lower than some state-of-the-art non-time-dense methods that utilize computationally intensive correlation volumes and focus solely on maximizing accuracy between two distant time points. Our current SMR module relies on implicit warp alignment for efficiency and time-density. Future work could explore hybrid approaches, potentially incorporating lightweight correlation features or attention mechanisms to enhance accuracy, especially for complex motions, without drastically increasing latency or computational cost.
  • Sparse Temporal Supervision: EVA-Flow achieves time-dense prediction but is still primarily supervised using ground truth optical flow provided at a much lower frame rate (e.g., 10Hz on DSEC). While our unsupervised RFWL metric helps validate intermediate flows, the network lacks direct, dense temporal supervision during training. Developing techniques for generating reliable time-continuous ground truth, perhaps through advanced interpolation or simulation, or exploring unsupervised/self-supervised learning objectives specifically designed for time-dense event flow could further enhance intermediate flow accuracy and reliability.
  • Event-Only Input: The current EVA-Flow framework operates solely on event data. While this highlights the richness of information within event streams, incorporating asynchronous image frames from the event camera (like Davis sensors) could provide complementary information, particularly in static scenes, low-texture areas, or during periods of low event activity. Designing efficient multi-modal fusion architectures that leverage both event dynamics and frame appearance is a promising direction for improving robustness and overall accuracy.
6 Conclusions
@@
In this paper, we propose EVA-Flow, a learning-based model for anytime flow estimation based on event cameras. Leveraging the Unified Voxel Grid representation, our network is able to estimate dense motion fields bin-by-bin between two temporal slices. By employing a novel Spatiotemporal Motion Refinement module for implicit warp alignment, the proposed EVA-Flow overcomes the low-temporal-frequency issue of event-based optical flow datasets, achieving high-speed and high-frequency flow estimation beyond low-frequency supervised signals. Our approach significantly outperforms other time-dense event flow methods in temporal capabilities and demonstrates exceptional computational efficiency (16.8 GMACs per prediction). It also sustains a competitive level of accuracy compared to non-time-dense techniques. Furthermore, we introduced the Rectified Flow Warp Loss (RFWL), an unsupervised metric for evaluating time-dense predictions, and showed that EVA-Flow’s prediction granularity can be flexibly adjusted during inference to suit varying scene dynamics. The combination of high temporal resolution, low latency, competitive accuracy, high computational efficiency, and adaptability makes EVA-Flow a highly practical and promising solution for real-time motion perception tasks on resource-aware platforms.
[…]
Conceptualization, Y.Y; methodology, Y.Y and H.S; software, Y.Y; validation, Y.Y and H.S; formal analysis, Y.Y and H.S; investigation, Y.Y; resources, X.Y, and Z.W; data curation, Y.Y, X.Y and Z.W; writing—original draft preparation, Y.Y; writing—review and editing, Y.Y, H.S, K.Y, L.S and K.W; visualization, Y.Y, H.S and L.S; supervision, K.Y, Y.W and K.W; project administration, K.W and Y.W; funding acquisition, K.W. All authors have read and agreed to the published version of the manuscript.
[…]
This research was funded by National Natural Science Foundation of China (Grant No. 12174341).
[…]
Not applicable.
[…]
Not applicable.
[…]
The data presented in this study are openly available in https://github.com/Yaozhuwa/EVA-Flow.
Acknowledgements.
[…]